log(x-1)^2=log(-5x-1)
Answer by stanbon(75887)

log(x-1)^2=log(-5x-1)
(x-1)^2 = -5x-1
x^2 - 2x + 1 = -5x-1
x^2 + 3x +2 = 0
(x+2)(x+1) = 0
x = -2 or x = -1
----------------------
Cheers,
Stan H.